OpenStudy (mtwizzle420):

Sarah can paint a small room in 45 minutes. Sam can paint that same room in 20 minutes. If they work together, how long will it take them to paint the room?

OpenStudy (mtwizzle420):

Im not asking for someone to do it for me but to help me out the whole way through. :)

OpenStudy (kropot72):

If Sarah works for 1 minute she can paint 1/45 of the room. If Sam works for 1 minute he can paint 1/20 of the room. Can you see what fraction of the room will be painted if they work together for 1 minute?

OpenStudy (mtwizzle420):

Thank you :) im going to review my notes really quick and try and set it up may I tag you in it once im ready? :)

OpenStudy (kropot72):

Please do :)

OpenStudy (mtwizzle420):

So I rewatched a live lesson and I have to use the Fraction Buster Method. Do you know what that is?

OpenStudy (kropot72):

If they work together for one minute the fraction of the room that will be painted is found as follows: \[\large \frac{1}{45}+\frac{1}{20}=\frac{4+9}{180}=\frac{13}{180}\]

OpenStudy (mtwizzle420):

whered the 4 and 9 come from?

OpenStudy (kropot72):

The addition of the two fractions 1/45 and 1/20 was performed by first finding the Lowest Common Multiple of 45 and 20, which is 180. Then both fractions are converted to 180ths and added. 1/45 = 4/180 and 1/20 = 9/180

OpenStudy (mtwizzle420):

is that the answer? or when its simplified is that the answer?

OpenStudy (mtwizzle420):

I got 20

OpenStudy (kropot72):

When both are working together for 1 minute they paint 13/180 of the room. Let t be the number of minutes needed for both working together to complete painting the room. Then we can write: \[\large t \times \frac{13}{180}=1\ .............(1)\] Now just solve equation (1) to find the value of t which is the time required in the question.

OpenStudy (mtwizzle420):

Okay hold on a sec :)

OpenStudy (mtwizzle420):

\[45\times\frac{ 13 }{ 180}=\frac{ 13 }{ 4 }\]

OpenStudy (mtwizzle420):

\[20\times \frac{ 13 }{ 180}=\frac{ 13 }{ 9 }

OpenStudy (mtwizzle420):

well for the last one I got 13 over 9

OpenStudy (kropot72):

\[\large t \times \frac{13}{180}=1\] Multiplying both sides by 180/13 we get: \[\large t=\frac{180}{13}=(you\ can\ calculate)\ minutes\]

OpenStudy (mtwizzle420):

Im very confused...

OpenStudy (kropot72):

Which step don't you understand?

OpenStudy (mtwizzle420):

So what I calculated was wrong , right?

OpenStudy (kropot72):

You don't seem to be doing what I asked.

OpenStudy (mtwizzle420):

And what is it you asked?

OpenStudy (kropot72):

When I asked you to solve: \[\large t \times \frac{13}{180}=1\ .............(1)\] you substituted 45 for t. That is not correct.

OpenStudy (kropot72):

To solve (1) just multiply both sides by 180/13 to get: \[\large t=\frac{180}{13}\]

OpenStudy (mtwizzle420):

OKay then.. what do I do for that step?

OpenStudy (kropot72):

Just divide 180 by 13 to get the time it would take to paint the room with Sarah and Sam working together.

OpenStudy (mtwizzle420):

Thank you !

OpenStudy (kropot72):

You're welcome :)